OpenStudy (anonymous):

The equation of the tangent line to f(x) = \sqrt{x} at x = 49 is y = . Using this, we find our approximation for \sqrt {49.3} is

OpenStudy (anonymous):

2 separate problems

OpenStudy (freckles):

Well first you need to find the tangent line of f(x)=sqrt(x) at x=49

OpenStudy (anonymous):

how do you do that

OpenStudy (freckles):

\[y-f(49)=f'(49)(x-49) \\ y=f'(49)(x-49)+f(49)\]

OpenStudy (freckles):

that is point slope form above

OpenStudy (freckles):

we just need to know the derivative of f and then plug in 49 into it we also need to know f(49) too

OpenStudy (anonymous):

I don't understand

OpenStudy (freckles):

which part?

OpenStudy (freckles):

finding f'?

OpenStudy (freckles):

ok i guess you got it since you closed it